# On the Provable Suboptimality of Momentum SGD in Nonstationary Stochastic Optimization

### v2

@@ 4.1 Strongly Convex Quadratic Minimizer
$\bm{\theta}_{t+1}^{\star}\;=\;\bm{\theta}_{t}^{\star}+\delta_{\mathrm{rw}}\,\frac{\mathbb{u}_{t}}{\|\mathbb{u}_{t}\|_{2}},\qquad\mathbb{u}_{t}\sim\mathcal{N}(0,I_{d}),$ (4.2)
which is a normalized random walk with step size $\delta_{\mathrm{rw}}$. This yields a cumulative drift scale $\sum_{t=1}^{T-1}\|\bm{\theta}_{t+1}^{\star}-\bm{\theta}_{t}^{\star}\|_{2}=(T-1)\delta_{\mathrm{rw}}$, matching the “drift budget” viewpoint common in tracking analyses. In Table˜1 we set $\delta_{\mathrm{rw}}=0.01$, but results were robust across varying levels of $\delta_{\mathrm{rw}}$. At each time step, the algorithm observes a noisy gradient
$G_{t}(\bm{\theta})=\nabla f_{t}(\bm{\theta})+\xi_{t}=\mu(\bm{\theta}-\bm{\theta}_{t}^{\star})+\xi_{t},$ (4.3)
where the noise $\xi_{t}\sim\mathcal{N}(0,\sigma^{2}I_{d})$ is an i.i.d. Gaussian vector with variance $\sigma^{2}$ in each coordinate. The performance metric is the squared tracking error: $e_{t}=\|\bm{\theta}_{t}-\bm{\theta}_{t}^{\star}\|^{2}.$ We report the averaged tracking error $\mathbb{E}[e_{t}]=\frac{1}{N}\sum_{i=1}^{N}\|\bm{\theta}_{t}^{(i)}-(\bm{\theta}_{t}^{\star})^{(i)}\|^{2}$ as a function of time $t$, where $(i)$ indexes independent runs. The experiment is averaged over 20 runs. These results were consistent across varying dimension values. In addition to Gaussian shifts, we also repeated the experiments with shifts drawn from a Student’s $t$ distribution, and observed qualitatively identical results.
4.2 Linear and Logistic Regression
Similar to the previous setup, to model non-stationarity we generate a time-varying population minimizer sequence $\{\bm{\theta}_{t}^{\star}\}_{t=1}^{T}$ via controlled drift. For linear and logistic regression, we set $\bm{\theta}_{t}^{\star}\in\mathbb{R}^{d}$ and evolve $\bm{\theta}_{t}^{\star}$ according to random walk in $\mathbb{R}^{d}$:
@@ 4.2 Linear and Logistic Regression
$\bm{\theta}_{t+1}^{\star}\;=\;\bm{\theta}_{t}^{\star}+\delta_{\mathrm{rw}}\,\frac{\mathbb{u}_{t}}{\|\mathbb{u}_{t}\|_{2}},\qquad\mathbb{u}_{t}\sim\mathcal{N}(0,I_{d}),$ (4.4)
which is a normalized random walk with step size $\delta_{\mathrm{rw}}$. Both of these tasks are run in two spectral regimes: well-conditioned ($\kappa=10$) and ill-conditioned ($\kappa=1000$). Conditioning is controlled by drawing covariates as $x=z\,\Sigma^{1/2}$ with $z\sim\mathcal{N}(0,I)$ and $\mathrm{cond}(\Sigma)=\kappa$, so that the smallest and largest eigenvalues satisfy $\mu=1$ and $L=\kappa$.
4.3 Teacher–student MLP drift.
For the teacher–student regression task, the population minimizer corresponds to the (time-varying) teacher network parameters. We also run this in both the well-conditioned and ill-conditioned regimes. Let $f_{\bm{\theta}}(x)$ be a two-layer ReLU MLP and let $\bm{\theta}_{t}^{\star}$ denote the teacher parameters at time $t$. We evolve $\bm{\theta}_{t}^{\star}$ analogously via either a continuous random-walk drift. Since parameter-space distances are not identifiable under hidden-unit permutation and ReLU scaling symmetries, we evaluate tracking primarily in prediction space i.e., the mean squared discrepancy between student and drifting teacher predictions on a fixed validation set.
4.4 Tables and Figures
@@
Table 1: Mean tracking error after 5000 iterations on a drifting strongly convex quadratic ($d=100$). Boldface indicates the best (lowest) method within each $(\beta,\sigma^{2},\gamma)$ setting. Across regimes, increasing $\beta$ markedly degrades the steady-state tracking of HB/NAG—especially at larger $\gamma$ and $\sigma^{2}$—highlighting inertia-induced lag under drift, while SGD remains comparatively robust.
$\gamma=0.01$
$\beta$ $\sigma^{2}$ SGD HB NAG
0.50 0.1 $1.036$ $0.342$ $0.349$
0.50 0.5 $1.235$ $0.728$ $0.745$
0.50 0.8 $1.305$ $0.961$ $1.019$
0.90 0.1 $1.029$ $0.497$ $0.453$
0.90 0.5 $1.230$ $2.358$ $2.247$
0.90 0.8 $1.466$ $3.899$ $3.721$
0.95 0.1 $1.039$ $1.051$ $0.808$
0.95 0.5 $1.263$ $5.144$ $4.114$
0.95 0.8 $1.351$ $7.696$ $6.801$
0.99 0.1 $1.036$ $4.837$ $2.619$
0.99 0.5 $1.231$ $23.669$ $13.096$
0.99 0.8 $1.403$ $38.802$ $21.038$
$\gamma=0.05$
$\beta$ $\sigma^{2}$ SGD HB NAG
0.50 0.1 $0.288$ $0.504$ $0.523$
0.50 0.5 $1.322$ $2.453$ $2.435$
0.50 0.8 $2.020$ $4.022$ $3.941$
0.90 0.1 $0.306$ $2.472$ $1.732$
0.90 0.5 $1.278$ $12.298$ $9.194$
0.90 0.8 $2.076$ $20.372$ $14.604$
0.95 0.1 $0.286$ $4.866$ $2.641$
0.95 0.5 $1.367$ $24.077$ $13.073$
0.95 0.8 $2.206$ $42.031$ $20.205$
0.99 0.1 $0.313$ $25.674$ $4.224$
0.99 0.5 $1.340$ $133.720$ $21.351$
0.99 0.8 $2.160$ $198.361$ $34.105$
$\gamma=0.10$
$\beta$ $\sigma^{2}$ SGD HB NAG
0.50 0.1 $0.528$ $1.054$ $0.914$
0.50 0.5 $2.626$ $5.273$ $4.848$
0.50 0.8 $4.112$ $8.505$ $7.813$
0.90 0.1 $0.525$ $5.172$ $2.596$
0.90 0.5 $2.540$ $27.411$ $13.433$
0.90 0.8 $4.286$ $40.666$ $21.376$
0.95 0.1 $0.534$ $9.981$ $3.331$
0.95 0.5 $2.663$ $51.361$ $17.378$
0.95 0.8 $4.036$ $79.822$ $27.169$
0.99 0.1 $0.545$ $54.250$ $4.414$
0.99 0.5 $2.715$ $243.978$ $24.099$
0.99 0.8 $4.110$ $401.372$ $38.317$
@@
Table 2: Benchmark summary across linear regression, logistic regression, and teacher–student MLP under streaming drift. Mean performance after $5000$ iterations for SGD, Heavy-Ball (HB), and Nesterov (NAG) under two conditioning regimes (well-conditioned: $\kappa=10$; ill-conditioned: $\kappa=1000$). Arrows indicate the preferred direction (lower is better for loss/tracking/val MSE; higher is better for accuracy). For the teacher–student MLP, tracking is computed in prediction space. We report the training settings $(\sigma^{2},\beta,\gamma)$ used in each regime (with $\gamma$ the base step size). Overall, SGD is the most robust across conditioning, while HB/NAG deteriorate sharply in ill-conditioned regimes—especially on tracking metrics—consistent with inertia amplifying lag under nonstationary drift.
@@
(a) $\gamma=0.1$, $\beta=0.9$, $\sigma^{2}=0.8$
(b) $\gamma=0.05$, $\beta=0.9$, $\sigma^{2}=0.5$
Figure 3: Tracking a drifting minimizer under strong convexity. Mean squared tracking error $\mathbb{E}\!\left[\|\bm{\theta}_{t}-\bm{\theta}_{t}^{\star}\|_{2}^{2}\right]$ versus time for SGD, Heavy-Ball (HB), and Nesterov (NAG) on a strongly convex quadratic with noisy gradients. Parameters $(\gamma,\beta,\sigma^{2})$ are shown above each panel with shaded bands denoting $\pm 1$ std over random seeds. Across both regimes, momentum methods (HB/NAG) suppress short-term noise but exhibit a substantially larger steady-state tracking error—consistent with inertia-induced lag when the minimizer drifts—whereas SGD tracks the minimizer more closely but with higher variability.

### v3

@@ F.1 General setup
$\bm{\theta}_{t+1}^{\star}=\bm{\theta}_{t}^{\star}+\delta_{\mathrm{rw}}\cdot\frac{\mathbf{u}_{t}}{\|\mathbf{u}_{t}\|_{2}},\qquad\mathbf{u}_{t}\sim\mathcal{N}(0,I_{d}),$
so that $\|\bm{\theta}_{t+1}^{\star}-\bm{\theta}_{t}^{\star}\|_{2}=\delta_{\mathrm{rw}}$ for each step. We set $\delta_{\mathrm{rw}}=0.01$ unless otherwise stated, and additionally report robustness to heavier-tailed Student-$t$ drift (Table˜3).
Initialization and training.
All optimizers are initialized at $\bm{\theta}_{0}=\bm{\theta}_{0}^{\star}$ with velocity buffer $\mathbf{v}_{0}=\mathbf{0}$, yielding zero initial tracking error to isolate the effect of drift. All experiments use mini-batch stochastic gradients with batch size $B=256$. The reported $\sigma^{2}$ values refer to label-noise variance for the teacher–student MLP). We use fixed base step sizes $\gamma$ for each task/regime (listed in the captions of Table˜3 and Table˜4) rather than per-method tuning. For the teacher–student MLP, teacher and student weights are initialized i.i.d. from $\mathcal{N}(0,0.04)$, and the student is warm-started to match the teacher at $t=0$.
Evaluation.
We report the squared tracking error $e_{t}=\|\bm{\theta}_{t}-\bm{\theta}_{t}^{\star}\|^{2}$ averaged over $N=20$ runs, with final performance measured at $T=5000$ iterations. Results were consistent across dimensions $d\in\{50,100,200\}$. The shaded regions in figures denote $\pm 1$ standard deviation.
F.2 Task-specific details
Strongly convex quadratic.
We consider the time-varying objective
$f_{t}(\bm{\theta})=\frac{\mu}{2}\|\bm{\theta}-\bm{\theta}_{t}^{\star}\|^{2},$ (238)
with $\mu=1$, so that $f_{t}$ is $1$-strongly convex with Lipschitz constant $L=\mu=1$ (condition number $\kappa=1$). The gradient is $\nabla f_{t}(\bm{\theta})=\mu(\bm{\theta}-\bm{\theta}_{t}^{\star})$. At each iteration, the algorithm receives a stochastic gradient
$G_{t}(\bm{\theta})=\nabla f_{t}(\bm{\theta})+\bm{\xi}_{t},\qquad\bm{\xi}_{t}\sim\mathcal{N}(0,\sigma^{2}I_{d}),$ (239)
with noise variance $\sigma^{2}\in\{0.1,0.5,0.8\}$. This task isolates the effect of gradient noise and momentum under drift in the simplest strongly convex setting.
Linear regression.
We consider streaming least-squares with drifting ground truth:
$y_{t}=\langle\mathbf{x}_{t},\bm{\theta}_{t}^{\star}\rangle+\epsilon_{t},\qquad\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2}),$ (240)
where covariates are drawn as $\mathbf{x}_{t}=\Sigma^{1/2}\mathbf{z}_{t}$ with $\mathbf{z}_{t}\sim\mathcal{N}(0,I_{d})$ and $d=50$. The covariance $\Sigma$ is constructed with eigenvalues log-spaced between $1$ and $\kappa$, yielding strong convexity parameter $\mu=1$ and smoothness $L=\kappa$. We test two conditioning regimes: well-conditioned ($\kappa=10$) and ill-conditioned ($\kappa=1000$).
Logistic regression.
We consider streaming binary classification with a drifting decision boundary:
$y_{t}\sim\mathrm{Bernoulli}\bigl(\varsigma(\langle\mathbf{x}_{t},\bm{\theta}_{t}^{\star}\rangle)\bigr),$ (241)
where $\varsigma(z)=1/(1+e^{-z})$ is the sigmoid function. Covariates are generated identically to linear regression, with the same conditioning regimes ($\kappa\in\{10,1000\}$) and dimension $d=50$. A small $\ell_{2}$ regularization term ($\lambda=10^{-3}$) is added to ensure strong convexity. We report tracking error $\|\bm{\theta}_{t}-\bm{\theta}_{t}^{\star}\|$, training loss, and classification accuracy evaluated on held-out samples.
Teacher–student MLP.
We consider a two-layer ReLU network of the form
$f_{\bm{\theta}}(\mathbf{x})=\mathbf{W}_{2}\,\mathrm{ReLU}(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1})+b_{2},$ (242)
with architecture and training details summarized in Table˜2. The teacher parameters $\bm{\theta}_{t}^{\star}=(\mathbf{W}_{1,t}^{\star},\mathbf{b}_{1,t}^{\star},\mathbf{W}_{2,t}^{\star},b_{2,t}^{\star})$ drift according to the same normalized random walk described above. Minimizers are generated as $y_{t}=f_{\bm{\theta}_{t}^{\star}}(\mathbf{x}_{t})+\epsilon_{t}$ with $\epsilon_{t}\sim\mathcal{N}(0,1)$.
Table 2: Teacher–student MLP architecture and training configuration.
Property Value
Architecture
     Input dimension ($d_{\mathrm{in}}$) 100
     Hidden dimension ($h$) 128
     Output dimension 1 (scalar)
     Activation ReLU
     Total parameters ($d_{\theta}$) 13,057
Initialization
     Weight distribution $\mathcal{N}(0,0.04)$
     Bias initialization $\mathcal{N}(0,0.04)$
     Student warm-start Matched to teacher at $t=0$
Training
     Batch size ($B$) 256
     Label noise ($\sigma$) 1.0
     Horizon ($T$) 5000
     Drift rate ($\delta_{\mathrm{rw}}$) 0.02
Evaluation
     Validation set size ($|\mathcal{V}|$) 2048
     Tracking metric Prediction-space MSE
     Seeds 20
Since parameter-space distances are not identifiable under the permutation and scaling symmetries of ReLU networks, we evaluate tracking in prediction space:
@@ F.2 Task-specific details
$e_{t}^{\mathrm{pred}}=\frac{1}{|\mathcal{V}|}\sum_{\mathbf{x}\in\mathcal{V}}\bigl\|f_{\bm{\theta}_{t}}(\mathbf{x})-f_{\bm{\theta}_{t}^{\star}}(\mathbf{x})\bigr\|^{2},$ (243)
where $\mathcal{V}$ is a fixed validation set held constant across all time steps. Conditioning is controlled via the input covariance $\Sigma$ as in the regression tasks.
F.3 Tables and figures
@@
Table 3: Mean tracking error after 5000 iterations on a drifting strongly convex quadratic ($d=100$) with Student-t distribution drift. Boldface indicates the best (lowest) method within each $(\beta,\sigma^{2},\gamma)$ setting.
$\gamma=0.01$ $\gamma=0.05$ $\gamma=0.10$
$\beta$ $\sigma^{2}$ SGD HB NAG SGD HB NAG SGD HB NAG
0.50 0.1 $0.056$ $0.103$ $0.100$ $0.256$ $0.509$ $0.459$ $0.573$ $1.079$ $0.942$
0.50 0.5 $0.245$ $0.498$ $0.509$ $1.286$ $2.597$ $2.491$ $2.464$ $5.206$ $4.750$
0.50 0.8 $0.383$ $0.792$ $0.760$ $2.043$ $4.151$ $3.611$ $4.226$ $8.010$ $7.785$
0.90 0.1 $0.054$ $0.501$ $0.471$ $0.257$ $2.397$ $1.831$ $0.532$ $5.015$ $2.689$
0.90 0.5 $0.241$ $2.585$ $2.282$ $1.232$ $13.022$ $8.099$ $2.607$ $24.212$ $13.258$
0.90 0.8 $0.400$ $3.947$ $3.637$ $1.999$ $21.154$ $14.204$ $4.236$ $40.316$ $21.831$
0.95 0.1 $0.051$ $1.015$ $0.830$ $0.255$ $5.350$ $2.514$ $0.507$ $9.995$ $3.378$
0.95 0.5 $0.237$ $4.930$ $4.062$ $1.256$ $23.270$ $12.708$ $2.703$ $50.767$ $16.846$
0.95 0.8 $0.404$ $7.718$ $6.624$ $2.101$ $40.283$ $20.850$ $3.947$ $79.354$ $28.313$
0.99 0.1 $0.052$ $4.767$ $2.433$ $0.268$ $26.093$ $4.490$ $0.520$ $51.724$ $4.656$
0.99 0.5 $0.247$ $26.309$ $12.036$ $1.226$ $131.242$ $21.845$ $2.673$ $245.091$ $23.187$
0.99 0.8 $0.391$ $40.600$ $21.424$ $1.968$ $210.841$ $33.517$ $4.143$ $405.148$ $38.306$
@@
Table 4: Benchmark summary across linear regression, logistic regression, and teacher–student MLP under streaming drift. Mean performance after $5000$ iterations for SGD, Heavy-Ball (HB), and Nesterov (NAG) under two conditioning regimes (well-conditioned: $\kappa=10$; ill-conditioned: $\kappa=1000$). Arrows indicate the preferred direction (lower is better for loss/tracking/val MSE; higher is better for accuracy). For the teacher–student MLP, tracking is computed in prediction space. We report the training settings $(\sigma^{2},\beta,\gamma)$ used in each regime (with $\gamma$ the base step size). Overall, SGD is the most robust across conditioning, while HB/NAG deteriorate sharply in ill-conditioned regimes which is consistent with inertia amplifying lag under nonstationary drift.